# Quantum Data Centers: Why Entanglement Changes Everything

### v1

@@ 4 Beyond Quantum Data Centers: Quantum Hubs
(a) Schematic representation of the considered physical quantum network architecture in [28]. The network comprises several QLANs. Within each QLAN, a orchestrator generates and distributes multipartite entangled states to a set of quantum nodes – referred to as clients – with a star-like topology. Inter-QLAN connectivity is enabled by point-to-point quantum channels, remarked in dark red, interconnecting different orchestrators.
@@
For the reasons highlighted in Sec. 3, each QLAN exhibits a centralized architecture, where the orchestrator handles the entanglement generation and distribution to the QLAN client nodes. This hierarchical design naturally extends to multi-Quantum Data Center networks, where inter-network physical connectivity is achieved through a mesh topology among the QLAN orchestrators, as represented in Fig. 9(a), while the inter-node communications leverage the entanglement-enabled connectivity. Indeed, this approach assures operational consistency across network scales, while respecting the constraints imposed by the technology maturity. As a consequence, from a quantum communication standpoint, the inter-node communication capabilities correspond to the generation of artificial links among network nodes belonging to different QLANS. Hence, as for intra-QLAN entanglement connectivity, the choice of the initial multipartite entangled state distributed in each QLAN is of paramount importance, for being able to adapt to different communication patterns [61, 62, 63].
In this context, [28] proposes four different prototypes to cater for different traffic patterns and connectivity requirements between QLANs: peer-to-peer, role delegation, clients hand-over and extranet. Each of these prototypes is associated with a different artificial topology designed to fulfill specific communication needs between two QLANs. Notably, these topologies are capable of dynamically creating multiple artificial links between distant nodes through local operations only (i.e., using free operations in the sense of quantum communication), without requiring new physical connections. This approach enhances the flexibility and adaptability of quantum networks while maintaining their resource efficiency. We strategically trade off these four traffic pattern prototypes depending on application requirements and network conditions, as follows.
@@
These four prototypes can all be realized between two different QLANs, as demonstrated in [28]. Moving forward, it becomes essential to consider how such a state can be engineered to: i) interconnect nodes belonging to different multiple QLANs, and ii) dynamically adapt to different inter-QLAN traffic demands. For example, how should we address the challenge of increasing the number of QLANs with maintaining the appropriate artificial topologies? How do the four pattern prototypes perform in a multi-QLAN architecture? To this end, we present the following preliminary ideas. By increasing the number of QLANs and nodes, we aim to keep the topology consistent while adjusting the entanglement resources accordingly. One possible direction is to explore the use of more scalable entangled resources — for example, extending from a bi-star state to an $n$-star graph state, as conceptually illustrated in Fig. 9(b)9(e) for the case of $n=3$. In this extended scenario, it remains an open question whether the artificial inter-QLAN topologies corresponding to primary archetypes — peer-to-peer, client hand-over, role delegation, and extranet — can be implemented using only local operations and measurements on such $n$-star entangled states. Investigating this possibility forms an important next step toward scalable and efficient multi-QLAN quantum network architectures.

### v2

@@ 4 Beyond Quantum Data Centers: Quantum Hubs
(a) Schematic representation of the considered physical quantum network architecture in [CheIllCac-24]. The network comprises several QLANs. Within each QLAN, a orchestrator generates and distributes multipartite entangled states to a set of quantum nodes – referred to as clients – with a star-like topology. Inter-QLAN connectivity is enabled by point-to-point quantum channels, remarked in dark red, interconnecting different orchestrators.
@@
For the reasons highlighted in Sec. 3, each QLAN exhibits a centralized architecture, where the orchestrator handles the entanglement generation and distribution to the QLAN client nodes. This hierarchical design naturally extends to multi-Quantum Data Center networks, where inter-network physical connectivity is achieved through a mesh topology among the QLAN orchestrators, as represented in Fig. 9(a), while the inter-node communications leverage the entanglement-enabled connectivity. Indeed, this approach assures operational consistency across network scales, while respecting the constraints imposed by the technology maturity. As a consequence, from a quantum communication standpoint, the inter-node communication capabilities correspond to the generation of artificial links among network nodes belonging to different QLANS. Hence, as for intra-QLAN entanglement connectivity, the choice of the initial multipartite entangled state distributed in each QLAN is of paramount importance, for being able to adapt to different communication patterns [CheIllCac-24-QCE, CheCacChe-23, WanRahRui-24].
In this context, [CheIllCac-24] proposes four different prototypes to cater for different traffic patterns and connectivity requirements between QLANs: peer-to-peer, role delegation, clients hand-over and extranet. Each of these prototypes is associated with a different artificial topology designed to fulfill specific communication needs between two QLANs. Notably, these topologies are capable of dynamically creating multiple artificial links between distant nodes through local operations only (i.e., using free operations in the sense of quantum communication), without requiring new physical connections. This approach enhances the flexibility and adaptability of quantum networks while maintaining their resource efficiency. We strategically trade off these four traffic pattern prototypes depending on application requirements and network conditions, as follows.
@@
These four prototypes can all be realized between two different QLANs, as demonstrated in [CheIllCac-24]. Moving forward, it becomes essential to consider how such a state can be engineered to: i) interconnect nodes belonging to different multiple QLANs, and ii) dynamically adapt to different inter-QLAN traffic demands. For example, how should we address the challenge of increasing the number of QLANs with maintaining the appropriate artificial topologies? How do the four pattern prototypes perform in a multi-QLAN architecture? To this end, we present the following preliminary ideas. By increasing the number of QLANs and nodes, we aim to keep the topology consistent while adjusting the entanglement resources accordingly. One possible direction is to explore the use of more scalable entangled resources — for example, extending from a bi-star state to an $n$-star graph state, as conceptually illustrated in Fig. 9(b)9(e) for the case of $n=3$. In this extended scenario, it remains an open question whether the artificial inter-QLAN topologies corresponding to primary archetypes — peer-to-peer, client hand-over, role delegation, and extranet — can be implemented using only local operations and measurements on such $n$-star entangled states. Investigating this possibility forms an important next step toward scalable and efficient multi-QLAN quantum network architectures.